Hi
Bill has 120 dollars more than James. Each week bill spends 25 dollars and James spends 22 dollars. They use up their money at the same time. How many weeks did it take bill to spend all his money.
thanks

Found 3 solutions by ankor@dixie-net.com, Theo, greenestamps:
Answer by ankor@dixie-net.com(22740)

Bill has 120 dollars more than James.
let j = amt of money James has
then
(j+120) = amt that Bill has
:
Each week bill spends 25 dollars and James spends 22 dollars.
They use up their money at the same time.
How many weeks did it take bill to spend all his money.
let w = no. of weeks to spend it all
:
Write an equation for each
Bill:
(j+120) - 25w = 0
Rearrange to
j - 25w = -120
James
j - 22w = 0
Use elimination on these two equation
j - 25w = -120
j - 22w = 0
--------------------Subtraction eliminates j, find w
0 - 3w = -120
w = -120/-3
w = 40 weeks both spend it all
:
:
To confirm this, find the actual amt each had.
j - 22(40) = 0
j - 880 = 0
j = $880 James had, and Bill had 120 more or $1000
880 - 40(22) = 0
1000 - 40(25) = 0

Answer by Theo(13342)

let b equal the amount of money that bill has.
let j equal the amount of money that james has.

bill has 120 dollars more than james.
that gets you b = j + 120.

bill spends 25 dollars a week and james spend 22 dollars a week.
they will both run out of money at the same time.

bill runs out when b - 25x = 0
james runs out when j - 22x = 0

since both expressions are equal to 0, they are equal to each other.
you get b - 25x = j - 22x
since b = j + 120, you get j + 120 - 25x = j - 22x.
subtract j from both sides of the equation and add 25x to both sides of the equation to get 120 = 3x
solve for x to get x = 40.

bill will run out of money in 40 weeks.
that's your solution.

james will also run out of money in 40 weeks.
bill will have spent 25 * 40 = 1000 dollars.
james will have spent 22 * 40 = 880 dollars.



Answer by greenestamps(13198)



Bill spends $3 more each week than James.

Bill starts with $120 more than James, and they run out of money at the same time.

The number of weeks it takes Bill to spend $120 more than James, spending $3 more than James each week, if 120/3=40.

ANSWER: 40 weeks